pooja195 (pooja195):

Solve the following system by substitution. 7x + 2y = 16 –21x – 6y = 24

pooja195 (pooja195):

I wanna try it out and just let me know if i go wrong :)

OpenStudy (anonymous):

Go for it. You can do it.

OpenStudy (anonymous):

Solve the following system by substitution. 7x + 2y = 16 -21x - 6y = 24 Start by isolating y

pooja195 (pooja195):

Wait O_O I forgot how to do these

OpenStudy (acxbox22):

do you have to do it by substitution or can u use any method? because elimination would be easier

OpenStudy (anonymous):

Isolate either variable (x or y) in either one of the equations. Choose whichever looks the easiest.

pooja195 (pooja195):

ok so if i subtracts 7x?

OpenStudy (anonymous):

In the first equation? Go for it.

pooja195 (pooja195):

2y=-7x+16 i dont know what to do after that? I mean would I divide by 2?

OpenStudy (anonymous):

That's right. Then you'll have an expression for y that you can substitute into the other equation. By the way, this problem will be a little trickier that the last one. But you can do it!

pooja195 (pooja195):

hahah thanks :) and i think i would get y = –( 7/2 )x + 8

OpenStudy (anonymous):

Don't forget the negative sign on the 7/2 x

pooja195 (pooja195):

right so y=-7/2x+8?

OpenStudy (anonymous):

Great. Now substitute this expression for y into the other equation. Simplify and solve for x.

pooja195 (pooja195):

–21(-7/2x+8) – 6y = 24

OpenStudy (anonymous):

Sorry. You substituted for x again. Remember the expression -7/2 x+8 is the expression for y and is substituted in place of y, not x

pooja195 (pooja195):

ugh lol i feel like im never gonna get this >_< –21x – 6(-7/2x+8)=24?

OpenStudy (anonymous):

Exactly. You got it. Now solve for x. I should warn you that there's a surprise coming up in this question, but it's a great learning opportunity.

OpenStudy (anonymous):

Don't forget the negative signs where they appear.

pooja195 (pooja195):

lol suprise hehe ok umm... 6(-7/2x+8)=24 im confused O_o

OpenStudy (anonymous):

Wait. you should now have\[-21x-6\left( -\frac{ 7 }{ 2 }x+8 \right)=24\]Multiply through by the 6 and simplify.

OpenStudy (anonymous):

See where we substituted for the y?

pooja195 (pooja195):

-21x-21x-48=24?

OpenStudy (anonymous):

That's it exactly. Now simplify and what do you get?

OpenStudy (anonymous):

Oops. There's an error with one of the signs. Double check you last result.

pooja195 (pooja195):

-21x+21x-48=24?

OpenStudy (anonymous):

That's better. Now simplify. What do you end up with?

pooja195 (pooja195):

-21x-21x=72??????? is that even possible O_O

OpenStudy (anonymous):

What is -21x+21x?

pooja195 (pooja195):

42?

OpenStudy (anonymous):

Just combine -21x and 21x. What do you get?

pooja195 (pooja195):

42 right?

pooja195 (pooja195):

42x?

OpenStudy (anonymous):

Nope\[-21x+21x=?\]

pooja195 (pooja195):

0

OpenStudy (anonymous):

Right on. So, you had\[-21x+21x-48=24\]Simplify the left hand side.

pooja195 (pooja195):

x=78?

OpenStudy (anonymous):

What is -21x + 21x again?

pooja195 (pooja195):

0

OpenStudy (anonymous):

Subtract 48 = ?

pooja195 (pooja195):

-48

OpenStudy (anonymous):

Right. So Everything you've done has come down to\[-48=24\]Is this possible?

pooja195 (pooja195):

no you would get a fraction =/

OpenStudy (anonymous):

SURPRISE!! Your work is correct. The problem is the question. There is no solution to that system of equations.

pooja195 (pooja195):

:OO all dis work for nothing

OpenStudy (anonymous):

I hear you. But take heart. What you've done is valuable. The process is always the same when solving by substitution. Solve one of the equations for one of the variables (doen's matter which...choose what looks easiest). Then substitute the resulting expression into the OTHER equation. You will then be able to solve that equation for one of the variables. Substitute that value into either of the original equations to solve for the other variable.

pooja195 (pooja195):

haha i get you thanks so so much <3 :D

OpenStudy (anonymous):

Feel like trying one that does have a solution? I could give you one.

pooja195 (pooja195):

haha sure :)

OpenStudy (anonymous):

Ok. Solve\[2x+3y=13\]\[4x-2y=2\]

pooja195 (pooja195):

ok so first I should add 2x right?

pooja195 (pooja195):

no subtract 2x

pooja195 (pooja195):

3y=-2x+14 divide by 3?

OpenStudy (anonymous):

For the first equation? Sure. But, there's a 13 in that equation, so you'll end up dealing ith fractions. Do you think it might be easier to work with the second equation?

pooja195 (pooja195):

yes so 4x−2y=2 add 4x -2y=-4x+2 divide by -2 y=-2x-1

OpenStudy (anonymous):

Terrific. The lesson hear is that it's worthwhile taking a second to look at each equation carefully and choosing which will be easier to work with. You can save yourself a lot of work. Now substitute this expression for y into the other equation.

pooja195 (pooja195):

7x + 2y = 16 7x+2(2x-1)=16 7x+4x-2=16 11x-2=16 11x=18 and im stuck >_<

OpenStudy (anonymous):

You're using the wrong equation. You should be substituting into 2x+3y=13.

pooja195 (pooja195):

oh i feel stupid now :( 2x+3(2x-1)=13. 2x+6x-3=13 8x-3=13 8x=16 divide by 8 x=2 :D

OpenStudy (anonymous):

Excellent. Know what to do next?

pooja195 (pooja195):

sub into the other problem right?

OpenStudy (anonymous):

Go for it.

pooja195 (pooja195):

7x + 2y = 16 7(2)+2y=16 14+2y=16 2y=2 y=1 :D

OpenStudy (anonymous):

Wrong equation again. Give it one more try. Either of the two equations I gave you.

pooja195 (pooja195):

I got confused from the previous problem 2x+3y=13 2(2)+3y=13

pooja195 (pooja195):

is that right so far?

OpenStudy (anonymous):

Yes.

pooja195 (pooja195):

4+3y=13 3y=9 y=3?

OpenStudy (anonymous):

Bravo! You must be exhausted! But you are now able to solve systems of two linear equations in two unknowns by substitution. Way to go, champ!

pooja195 (pooja195):

Hahaha thank you so much I actually get this I appreciate your efforts your awesome <3

OpenStudy (anonymous):

Any time.